Show that the vector →{v}=(yz-1)i-z(1+x+z)j+y(1+x+2z)k is conservative and find it scalar potential function.


1
Expert's answer
2021-08-12T07:27:30-0400

Given:

"{\bf v}=(yz-1){\bf i}+z(1+x+z){\bf j}+y(1+x+2z){\bf k}"


For the conservative vector field

"\rm curl {\bf v}=\bf 0"

We have

"\rm curl {\bf v}=\begin{vmatrix}
 {\bf i} & {\bf j} & {\bf k}\\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z}\\
(yz-1)&z(1+x+z)& y(1+x+2z)
\end{vmatrix}"

"=(1+x+2z-(1+x+2z)){\bf i}"

"+(y-y){\bf j}+(z-z){\bf k}=\bf 0"

Therefore, the given vector field is conservative.

We can put

"{\bf v}=\nabla\psi"

where

"\psi=yz(1+x+z)-x"